Jim Keys holds a $200,000 portfolio consisting of the following stocks:

Stock

Investment

Beta

Alpha

$50,000 0.50

Beta

$50,000 0.80

Gamma

$50,000 1.00

Delta

$50,000 1.30


What is the portfolio's beta?
Answer #1
Total Portfolio value = Value of Alpha + Value of Beta + Value of Gamma + Value of Delta
=50000+50000+50000+50000
=200000
Weight of Alpha = Value of Alpha/Total Portfolio Value
= 50000/200000
=0.25
Weight of Beta = Value of Beta/Total Portfolio Value
= 50000/200000
=0.25
Weight of Gamma = Value of Gamma/Total Portfolio Value
= 50000/200000
=0.25
Weight of Delta = Value of Delta/Total Portfolio Value
= 50000/200000
=0.25
Beta of Portfolio = Weight of Alpha*Beta of Alpha+Weight of Beta*Beta of Beta+Weight of Gamma*Beta of Gamma+Weight of Delta*Beta of Delta
Beta of Portfolio = 0.5*0.25+0.8*0.25+1*0.25+1.3*0.25
Beta of Portfolio = 0.9